OpenStudy (precal):

Consider the following piece-wise defined function. Part I) If a=-3 and b=4, will f(x) be continuous at x=2? Part II) If a=-3 and b=4, will f(x) be differentiable at x=2? Part III) For what vales of a and b will f(x) be both continuous and differentiable at x=2?

OpenStudy (precal):

ax^2+bx+2, x is less than and equal to 2 ax+b, x >2

OpenStudy (precal):

Part 1) I sub a=-3 b=4 and x=2 into f(x) and determine -2=-2 so therefore f(x) is continuous at x=2 Part 2) I sub a=-3 b=4 and x=2 into f'(x) and determine that the limit of f'(x) DNE at x=2

OpenStudy (precal):

or I should say the limit as x approaches 2 of f'(x) DNE

OpenStudy (anonymous):

what? solve for a and b.

OpenStudy (precal):

Part 3) is where I got stuck I took f(x) and sub x=2 and solve it, now I think I should have taken the derivative of the piecewise and sub x=2 and then solve it

OpenStudy (precal):

Yes part 3) wants the a and b value that will make it both continuous and differentiable

OpenStudy (anonymous):

Okay, so you found f'(x)?

OpenStudy (precal):

doing it right now

OpenStudy (precal):

I think it is a= -(1/3)b

OpenStudy (precal):

not sure that is correct

OpenStudy (anonymous):

What did you get for f'(x)?

zepdrix (zepdrix):

Hmm that's the same thing I'm coming up with.

OpenStudy (precal):

2ax+b for x<2 a for x>2 then I set that equal to each other 2ax+b=a 2a(2)+b=a 4a+b=a b=-3a b/-3=a

OpenStudy (precal):

so did I take the right approach I know that differentiablity implies continuity I at not sure I spelled that correctly

zepdrix (zepdrix):

You can solve for a and b a little further though. You should get a system of two equations by equating the left and right limits of f(x), and also equating the left and right limits of f'(x)

OpenStudy (anonymous):

The second equation for a and b comes form f(2).

OpenStudy (anonymous):

sub -b/3 into a.

zepdrix (zepdrix):

This gives you an equation:\[\Large\rm \lim_{x\to2^-}ax^2+bx+2\quad =\quad \lim_{x\to2^+}ax+b\]And so does this:\[\Large\rm \lim_{x\to2^-}2ax+b\quad=\quad \lim_{x\to2^+}a\]

OpenStudy (precal):

ok not sure what to do, I am getting a=0 b=0

OpenStudy (precal):

I don't think that is correct

zepdrix (zepdrix):

The system: \[\Large\rm 4a+2b+2=2a+b\]\[\Large\rm 4a+b=a\]Hmm I came up with a=2, I haven't checked b yet..

OpenStudy (precal):

sorry did not think to go back to f(x)

OpenStudy (precal):

yes a=2 and b=-6

zepdrix (zepdrix):

cool c:

OpenStudy (precal):

Thanks :)